The frequency of vibration of a mass m suspended from a spring constant k us given by the relation F=Cm^xK^y is the dimensionless constant determine x and y

1
Expert's answer
2022-02-02T05:14:48-0500

The frequency of vibration of a mass "m" suspended from a spring constant "k" us given by the relation

"f\sim m^xk^y"

Let's apply the dimensional analysis. We get

"[f]=T^{-1},\; [m]=M,\; [k]=MT^{-2}"

"T^{-1}=M^x(MT^{-2})^y"

Hence

"x+y=0,\\-2y=-1."

Finally

"y=1/2,\; x=-1/2"

"f\sim\sqrt{\frac{k}{m}}"